# Inverse 3D microscopy rendering for cell shape inference with active mesh

### v2

@@ Abstract
Traditional methods for biological shape inference, such as deep learning (DL) and active contour models, face limitations in 3D. DL requires large labeled datasets, which are difficult to obtain, while active contour models rely on fine-tuned hyperparameters for intensity attraction and regularization. We introduce deltaMic, a novel 3D differentiable renderer for fluorescence microscopy. By leveraging differentiable Fourier-space convolution, deltaMic accurately models the image formation process, integrating a parameterized microscope point spread function and a mesh-based object representation. Unlike DL-based segmentation, it directly optimizes shape and microscopy parameters to fit real microscopy data, removing the need for large datasets or heuristic priors. To enhance efficiency, we develop a GPU-accelerated Fourier transform for triangle meshes, significantly improving speed. We demonstrate deltaMic’s ability to reconstruct cellular shapes from synthetic and real microscopy images, providing a robust tool for 3D segmentation and biophysical modeling. This work bridges physics-based rendering with modern optimization techniques, offering a new paradigm for microscopy image analysis and inverse biophysical modeling.
@@ 1 Introduction
Figure 1: Graphical abstract of shape and microscope PSF inference with our differentiable fluorescence microscopy 3D renderer.
@@
Fluorescence microscopy [44] is the most widely used technique for imaging biological structures. However, extracting quantitative information from 3D fluorescence images remains a major bottleneck, limiting the development of new biological shape analysis methods and models. In fluorescence microscopy, biological samples are labeled with fluorescent dyes (fluorophores) that bind to specific structures of interest—such as lipid membranes, cytoskeletal networks, or organelles—making them bright while keeping the background dark. Except for unavoidable camera or biological noise, 3D fluorescent images are typically sparse and composed of well-structured objects such as points, filaments, surfaces, or bulk solids.
@@
Leveraging this inherent structure, deep-learning-based tools tailored for fluorescence imaging [4], particularly convolutional neural networks (CNNs) [65, 12], have enabled significant advances in automating challenging tasks such as image restoration [76], instance segmentation [67, 71], and feature encoding via self-supervised learning [36]. Despite these achievements, the expressivity of neural networks also presents challenges in 3D biological imaging: 1.Data requirements: 3D CNNs and alternative architectures (e.g., Vision Transformers) require large annotated datasets, which are difficult to produce and label accurately in 3D. While foundation models for biomedical image analysis [47] show promise, commonly used deep-learning (DL) models remain highly dependent on specific microscope imaging modalities, requiring extensive retraining when imaging conditions change. 2.Lack of interpretability and physical constraints: Many DL models used in laboratories [71] lack interpretability and rarely incorporate prior knowledge about optical image formation, limiting their reliability and scientific insight. 3.Geometric analysis and shape tracking: CNNs typically produce segmentation masks, which are suboptimal for measuring geometric features like curvatures or angles [28]. Additionally, timelapse microscopy requires object shape tracking [74], yet most segmentation models operate frame by frame, necessitating a separate shape-matching step as post-processing. 4.Risk of hallucinations: DL models can generate artifacts or structures that do not correspond to real biological features—an issue particularly problematic for scientific research and medical applications, where accuracy is critical.
@@
In response to these challenges, a growing trend in microscopy image analysis favors methods that incorporate stronger prior knowledge about optical physics [3, 41, 69, 2]. In contrast to DL, active contour methods (also called active snakes or active meshes, depending on the parameterization), which have been developed for decades [11, 31], directly evolve a shape representation to minimize an objective function that aligns with the underlying image [15, 48, 16, 73, 68]. However, these methods rarely incorporate the physical principles of fluorescence image formation directly, instead relying on user-controlled regularization terms to smooth output shapes. While largely supplanted by DL due to efficiency and speed, they retain advantages such as the ability to track objects over time and to produce shape representations like meshes or level-sets, which integrate seamlessly with modeling frameworks.
@@
Meanwhile, in computer vision, the development of automatic differentiation libraries such as PyTorch [62] and JAX [20] has accelerated the adoption of differentiable models. These tools have been pivotal in inverse differentiable rendering [32, 58] and mesh-based shape inference methods from images [70]. Although inverse differentiable rendering has demonstrated its effectiveness in computer vision, it remains largely unexplored in fluorescence image analysis. Here, we introduce a novel paradigm that leverages the optical properties of fluorescence microscopy to approximate the image formation process from biological object shapes parameterized with a mesh. We present deltaMic, a differentiable model that renders 3D fluorescence microscopy images from surface meshes, which can be directly compared to real biological images using a voxel-based weighted L2 norm. Implemented in PyTorch for automatic differentiation, our approach efficiently optimizes this loss by computing gradients with respect to both the biological object’s shape and the microscope’s optical properties. As a result, our method enables precise biological shape inference while simultaneously emulating the optics of a fluorescence microscope, eliminating the need for per-sample hyperparameter tuning required in traditional active contour methods.
@@
$\bullet$ We introduce a simplified model of 3D fluorescence microscopy, combining a mesh-based object representation with a parameterized point spread function (PSF).
$\bullet$ We present a differentiable Fourier transform for triangle surface meshes and optimized GPU implementation.
$\bullet$ We demonstrate that our differentiable microscopy renderer can accurately reconstruct 3D cellular shapes from both synthetic and real 3D microscopy images without requiring additional shape-regularization terms.
@@ 2.1 Fluorescence microscopy modeling
In fluorescence microscopy, a laser excites fluorescent dyes bound to the biological structure of interest at a specific wavelength, prompting them to emit fluorescence at a longer wavelength. This emitted light passes through the microscope optics and is detected by a CMOS camera, producing a 2D image. In confocal or light-sheet microscopy, the focal plane is incrementally adjusted to generate a 3D volumetric image composed of optical sections of the sample. However, the resolution of these images is fundamentally limited by the diffraction of light through the microscope optical elements [29]. When the imaging system’s response is translation-invariant in 3D space, it can be characterized by a single point spread function (PSF), which describes how a point source is imaged. Various physics-based PSF models have been developed [21, 38, 23, 24, 1], primarily for deconvolution applications [66]. The most straightforward method to determine a PSF is to generate experimental microscopy images of micrometric fluorescent beads that approximate point sources. PSF models can then be assessed by fitting them to experimentally acquired PSF images. In our approach, we approximate the PSF using the simplest possible model: a Gaussian kernel, that is fully characterized by its covariance matrix.
Knowing the PSF also enables the creation of artificial microscopy images, a topic explored in several studies. Synthetic fluorescence image generation can be broadly categorized into two approaches: models that aim to replicate the real image-formation process [14] and methods leveraging texture synthesis [51, 77] or generative deep learning [26, 17] to produce photorealistic images. These synthetic images are primarily used for evaluating image-analysis algorithms [63] or generating annotated datasets for training deep learning models [54]. In [14], the fluorophore distribution is represented by a Boolean mask indicating fluorophore presence, which is then convolved with a user-provided PSF before adding camera noise. We build upon and extend this approach by defining a fluorophore density over each simplex of the mesh rather than using discrete point sources or a mask. This avoids the need to densely populate biological shape meshes with millions of points, significantly reducing computational costs.
@@ 2.2 Instance segmentation of fluorescent images
Biological image segmentation has evolved similarly to computer vision [52], progressing from traditional methods such as thresholding, watershed transform [5, 6, 18], and active contours [31, 11, 48, 16] to graph-cut optimization [7], before the emergence of deep learning (DL)-based pipelines [65, 12]. In 2D, the most effective segmentation methods leverage large, diverse annotated datasets to train convolutional neural networks (CNNs) [61, 37, 22] or, more recently, vision Transformers [47]. These models predict instance masks robustly and generically, independent of specific biological samples or imaging modalities. However, replicating this success in 3D has proven challenging, demonstrating that DL-based volumetric image analysis is not merely an extension of 2D approaches. 3D images, or z-stacks, are obtained by sequentially acquiring 2D images at different focal depths along the optical axis. This process introduces significant anisotropy along the z-direction, which varies across imaging conditions, resulting in heterogeneous datasets that complicate CNN training and generalization. Additionally, the large size of modern 3D biological images—ranging from $512^{3}$ voxels in confocal microscopy to $2048^{3}$ in light-sheet imaging—poses substantial memory constraints for current GPU architectures. Another major challenge is labeling volumetric data. This task demands expert knowledge of the biological system, strong spatial visualization skills to ensure consistent annotations across planes, and efficient labeling software. Moreover, transitioning from 2D to 3D significantly increases both the number of voxels to annotate and the complexity of the task
A significant portion of classical image segmentation relies on minimizing an energy functional $\mathcal{E}(\Lambda,m)$, which models the distance between a target shape $\Lambda$ and distinctive image features m [55]. These features may include sharp intensity gradients (using edge detectors [50]) or regions of distinct intensities [11]. Shape representation is typically handled using either implicit level sets or explicit meshes, with the optimal shape $\Lambda^{*}$ obtained through gradient-based optimization. The negative of the gradient $\frac{\partial\mathcal{E}(\Lambda,m)}{\partial\Lambda}$ acts as a force guiding the contour toward the target shape, inspiring the terms active snakes, contours [31, 48, 73], or meshes [16, 68]. However, these methods often require user-defined regularization terms (akin to tension or bending energies) to enforce smoothness and prevent sharp artifacts. Fluorescence microscopy images of slender biological structures, such as membranes, pose a challenge for traditional region- or edge-based energies, as these structures appear as thin, high-intensity regions against a dark background. Alternative approaches define forces based on proximity to local intensity maxima [48, 72] or introduce energy functionals that compare real microscopy images to synthetic images generated from a mesh [25, 57]. The synthetic image is typically produced by convolving a binary mask derived from the mesh with a known PSF. Our approach builds on these ideas but introduces a more rigorous and versatile 3D rendering framework. Unlike previous methods, we jointly optimize the mesh and PSF parameters within a differentiable pipeline, eliminating the need for explicit shape regularization while achieving robust and physically consistent reconstructions.
@@ 2.3 Differentiable rendering
Rendering 2D or 3D geometrical shapes into rasterized 2D images is a fundamental topic in computer graphics. Recent advances have made rendering frameworks increasingly differentiable, enabling inverse-rendering problems where scene parameters (e.g., shapes, textures, and material properties) can be learned directly from raster images [43, 59, 33]. However, standard rasterization pipelines are not inherently differentiable, as discontinuities and occlusions introduce non-smooth gradients. To address this, various smoothing strategies have been developed for differentiable rasterization [64, 46, 39, 34]. Our fluorescence microscopy emulator functions as a 3D differentiable renderer. Similar to differentiable rendering in computer graphics, it relies on mathematically differentiable operations, here implemented in PyTorch. In contrast to computer graphics, the image produced lives in 3D space. The image formation process is naturally smoothed by convolving the Fourier transform (FT) of the mesh with a PSF, akin to the smoothing strategies used in differentiable rasterization [46]. Given that a triangle mesh typically consists of thousands of vertices and a PSF can involve hundreds of parameters, optimizing such models requires efficient differentiation techniques. Assuming all model components consist of differentiable operations, reverse-mode differentiation (backpropagation) efficiently computes gradients via the chain rule. The increasing demand for efficiency and scalability has led to the development of GPU-accelerated automatic differentiation libraries [62, 20]. We leverage PyTorch’s [62] optimized backpropagation framework and implement our mesh FT as a differentiable function within this pipeline.
3 3D fluorescence microscopy image rendering
Images are represented as intensity maps from $[0,1]^{3}$ to $\mathbb{R}$ without loss of generality, as any non-cubic image can be linearly rescaled to fit within $[0,1]^{3}$. In the following, we introduce approximated models for both the imaging system and the geometry of biological samples, which are then integrated to construct our differentiable renderer.
Figure 2: Principle of 3D fluorescence microscopy rendering: A continuous fluorophore density defined on the cell surface $\Lambda$ is convolved with the imaging system’s PSF to generate a 3D fluorescence image.
@@ 3.1 Translation-invariant rendering model
We assume that the PSF $h$ of the fluorescence microscope is translation-invariant. Under this assumption, a smooth image $u_{\alpha}:[0,1]^{3}\to\mathbb{R}$ is obtained by convolving the fluorophore density $u_{\Lambda}:[0,1]^{3}\to\mathbb{R}$ with the point spread function (PSF) kernel $h:[0,1]^{3}\to\mathbb{R}$ :
$u_{\alpha}(\textbf{p})=(u_{\rho}*h)(\textbf{p})=\int_{[0,1]^{3}}u_{\Lambda}(\textbf{x})h(\textbf{p}-\textbf{x})\mathrm{d}^{3}\textbf{x}.$ (1)
To circumvent the direct computation of these integrals, we perform the convolution in Fourier space by applying an element-wise multiplication of the Fourier transform (FT- $\hat{u}_{\Lambda}$ and $\hat{h}$, yielding: $\hat{u}_{\alpha}=\hat{u}_{\Lambda}\cdot\hat{h}.$
With this image formation model, one can start from a first guess $\left(u_{\Lambda}^{0},h^{0}\right)$, and minimize the distance between the rendered image $u_{\alpha}^{0}$ and a real microscopy image $m$. Doing so allows to infer both the fluorophore density of the observed biological sample and the PSF of the imaging system, $\left(u_{\Lambda}^{*},h^{*}\right)$. However, to learn meaningful shape representations, both of these functions have to be parametrized, using prior knowledge on the shape considered and its topology, as well as on the PSF.
With this image formation model, one can start with an initial estimate $\left(u_{\Lambda}^{0},h^{0}\right)$ and minimize the difference between the rendered image $u_{\alpha}^{0}$ and a real microscopy image $m$. This optimization process enables the inference of both the fluorophore density of the biological sample and the PSF of the imaging system, yielding $\left(u_{\Lambda}^{*},h^{*}\right)$. However, to obtain meaningful shape representations, both functions must be parameterized, incorporating prior knowledge about the object’s shape, topology, and the PSF.

### v3

@@ Abstract
Traditional methods for biological shape inference, such as deep learning (DL) and active contour models, face important limitations in 3D. DL approaches require large annotated datasets, which are often impractical to obtain, while active contour methods depend on carefully tuned heuristics for intensity attraction and shape regularization. We introduce deltaMic, a novel differentiable 3D renderer for fluorescence microscopy that formulates shape inference as an inverse problem. By leveraging differentiable convolutions, deltaMic simulates the image formation process, integrating a parameterized point spread function (PSF) with a triangle mesh-based representation of biological structures. Unlike DL- or contour-based segmentation, deltaMic directly optimizes both shape and optical parameters to align synthetic and real microscopy images, removing the need for large datasets or sample-specific fine-tuning. To ensure scalability, we implement a GPU-accelerated Fourier transform for triangle meshes along with narrow-band spectral filtering. We show that deltaMic accurately reconstructs cell geometries from both synthetic and diverse experimental 3D microscopy data, while remaining robust to noise and initialization. This establishes a new physics-informed framework for biophysical image analysis and inverse modeling.
@@ 1 Introduction
Figure 1: Graphical summary: joint inference of cell geometry and optical PSF using our differentiable 3D fluorescence microscopy.
@@
Fluorescence microscopy [51] is the most widely used technique for imaging biological structures. However, extracting quantitative information from 3D fluorescence images remains a major bottleneck, limiting the development of new biological shape analysis methods and models. In fluorescence microscopy, biological samples are labeled with fluorescent dyes (fluorophores) that bind to specific structures of interest—such as lipid membranes, cytoskeletal networks, or organelles—making them bright while keeping the background dark. Except for unavoidable camera noise and biological heterogeneities, 3D fluorescent images are typically sparse and composed of well-structured objects such as points, filaments, surfaces, or bulk solids.
@@
Leveraging this inherent structure, tools based on deep-learning (DL) tailored for fluorescence imaging [5], particularly convolutional neural networks (CNNs) [72, 13], have enabled significant advances in automating challenging tasks such as image restoration [83], instance segmentation [74, 78], and feature encoding via self-supervised learning [44]. Despite these achievements, the expressivity of neural networks also presents challenges in 3D biological imaging: (1) Data requirements: 3D CNNs or newer architectures (e.g., Vision Transformers [19]) require large annotated datasets, which are difficult to produce and label accurately in 3D. While foundation models for biomedical image analysis show promise [54], commonly used DL models remain highly dependent on specific microscope imaging modalities, requiring extensive retraining when imaging conditions change. (2) Lack of interpretability and physical constraints: widely-used DL segmentation models [78] lack interpretability and rarely incorporate prior knowledge about optical image formation, limiting their scientific insight. (3) Geometric analysis: CNNs typically produce segmentation masks, which are suboptimal for measuring geometric features like curvatures or angles [34]. Additionally, timelapse microscopy requires object shape tracking [81], yet most segmentation models operate frame by frame, necessitating a separate shape-matching step as post-processing. (4) Risk of hallucinations: DL models can generate artifacts or structures that do not correspond to real biological features—an issue for scientific research and medical applications, where accuracy is critical.
@@
In response to these challenges, a growing trend in microscopy image analysis favors methods that incorporate stronger prior knowledge of optical physics [4, 48, 76, 3]. In contrast to DL, active contour methods—also known as active snakes or meshes—directly evolve a shape representation to minimize an objective function aligned with image features [12, 38, 20, 55, 21, 80, 75]. Although being developed for decades, these methods have seldom incorporated the physical principles of fluorescence image formation, instead relying on user-defined regularization terms to enforce smoothness. While DL approaches recently supplanted them due to superior speed and automation, active contour methods retain valuable advantages: they can track objects over time and produce explicit shape representations (e.g., meshes, level sets) that ease geometric analysis and integrate naturally with downstream modeling frameworks.
@@
Meanwhile, the emergence of libraries such as PyTorch [69] and JAX [26], which offer native automatic differentiation, has accelerated the adoption of differentiable models in computer vision (CV), notably in inverse rendering [41, 65] and mesh-based shape inference [77]. Although inverse rendering has proven effective in CV, it remains largely unexplored in fluorescence biological image analysis. Here, we introduce a novel paradigm that leverages the fundamental optics of fluorescence microscopy to approximate the 3D image formation process from biological object shapes. We present deltaMic, a differentiable model that renders 3D fluorescence microscopy images from surface meshes, which can be directly compared to real biological images using a weighted voxel-based $L^{2}$ norm. Implemented in PyTorch for automatic differentiation, our approach efficiently optimizes this loss by computing gradients with respect to both the biological object’s shape and the microscope’s optical properties. As a result, our method enables precise biological shape inference while simultaneously emulating the microscope’s optics, eliminating the need for per-sample hyperparameter tuning required in traditional active contour methods.
@@
$\bullet$ We introduce a simplified model of 3D fluorescence microscopy generation, combining a mesh-based object representation with a parameterized point spread function (PSF).
$\bullet$ We implement a differentiable Fourier transform of triangle surface meshes parallelized on GPU.
$\bullet$ We demonstrate that our differentiable microscopy renderer can accurately infer 3D cellular shapes from both synthetic and real 3D microscopy images without requiring additional shape-regularization terms.
$\bullet$ We provide a PyTorch implementation on GitHub [35].
@@ 2.1 Fluorescence microscopy modeling
In fluorescence microscopy, a laser excites fluorescent dyes bound to biological structures at a specific wavelength, causing them to emit light at a longer wavelength. The emitted fluorescence passes through the optics and is captured by a CMOS camera, producing a 2D image. In modalities such as confocal or light-sheet microscopy, the focal plane is incrementally shifted to acquire a 3D volumetric image composed of optical sections [42]. However, resolution is fundamentally limited by the diffraction of light through the optical system [36]. When the imaging system’s response is approximately translation-invariant in 3D, it can be described by a point spread function (PSF), which characterizes how a point source is imaged. Various physics-based PSF models have been proposed [27, 45, 29, 30, 2], mainly for deconvolution purposes [73]. The simplest way to determine a PSF is to image sub-micrometer fluorescent beads approximating point sources. The model is then fitted to these experimental images. In this work, we approximate the PSF using a basic Gaussian kernel for simplicity; however, the framework could also incorporate more accurate experimental or physics-based models [1].
Knowing the PSF also enables synthetic image generation, explored in several studies. These approaches either replicate the real image-formation process [18] or use texture synthesis [58, 84] or generative DL [32, 23] to produce realistic images. Such images are primarily used to evaluate image-analysis algorithms [70] or generate annotated datasets for DL training [61]. In [18], the fluorophore distribution is modeled as a Boolean mask convolved with a user-provided PSF and noise. We extend this approach by defining a fluorophore density over each mesh simplex rather than using discrete sources or a mask. This avoids densely populating biological meshes with millions of points and significantly reduces computation.
@@ 2.2 Instance segmentation of fluorescent images
Biological image segmentation has evolved in parallel with computer vision [59], from traditional methods such as thresholding, watershed [6, 7, 24], and active contours [38, 12, 55, 21], to graph-cut optimization [8], and more recently DL-based pipelines [72, 13]. In 2D, the most effective methods leverage large annotated datasets to train CNNs [68, 14, 28] or, more recently, vision Transformers [54]. These models predict instance masks robustly across samples and imaging modalities. However, extending this success to 3D remains challenging.
3D images (z-stacks) are acquired by capturing 2D slices at varying depths, introducing anisotropy along the z-axis that varies with imaging conditions, yielding heterogeneous datasets that hinder CNN training and generalization. Modern 3D images—ranging from $512^{3}$ voxels in confocal to $2048^{3}$ in light-sheet microscopy—also impose substantial memory constraints on GPUs. Labeling 3D data poses major challenges. It requires expert biological knowledge, strong spatial skills to maintain cross-plane consistency, and suitable annotation tools. The shift from 2D to 3D increases both the number of voxels and annotation complexity.
Classical contour-based segmentation techniques often minimize an energy functional $\mathcal{E}(\Lambda,m)$ modeling the distance between a shape $\Lambda$ and image features $m$ [62], such as intensity gradients [57] or homogeneous regions [12]. Shape representations use level sets or meshes, and the optimal shape $\Lambda^{*}$ is found via gradient-based optimization. The negative gradient $\frac{\partial\mathcal{E}(\Lambda,m)}{\partial\Lambda}$ acts as a force guiding the contour—hence the terms active snakes, contours [38, 55, 80], or meshes [21, 75]. These methods typically require user-defined regularization (e.g., penalization terms akin to tension or bending energies) to maintain smoothness and avoid artifacts.
Fluorescence images of slender structures like membranes are challenging for region- or edge-based energies, as such structures appear as thin, high-intensity features against a dark background. Alternatives use forces based on proximity to local intensity maxima [55, 79] or compare real and synthetic images generated from meshes [31, 64], where the synthetic image is produced by convolving a binary mesh mask with a known PSF. Our approach builds on this idea but introduces a more rigorous 3D rendering framework and direct variational formulation. We jointly optimize the mesh and PSF parameters in a differentiable pipeline, eliminating the need for explicit shape regularization while maintaining robust and physically grounded reconstructions.
@@ 2.3 Differentiable rendering
Rendering 2D or 3D geometrical shapes into rasterized images is a core problem in computer graphics. Recent advances have enabled differentiable rendering, allowing inverse problems where scene parameters (e.g., shapes, textures, materials) are learned from raster images [50, 66, 40]. Traditional rasterization is non-differentiable due to occlusions and discontinuities, prompting development of smoothing strategies [71, 53, 46, 39].
Our fluorescence microscopy emulator functions as a 3D differentiable renderer. Like in graphics, it relies on mathematically differentiable operations. Unlike graphics, the resulting image lives in 3D space. Image formation is smoothed by convolving the mesh with a PSF, similarly to smoothing in differentiable rasterization [53]. Since a triangle mesh may include thousands of vertices and a PSF hundreds of parameters, optimization demands efficient differentiation. Assuming all components are differentiable, reverse-mode differentiation (backpropagation) computes gradients via the chain rule. Growing demands for scalability have led to GPU-accelerated automatic differentiation libraries [69, 26]. We leverage PyTorch’s backpropagation engine and implement our mesh Fourier Transform (FT) as a differentiable function within it.
Figure 2: DeltaMic pipeline overview: The microscopy renderer generates a 3D synthetic fluorescence image from a triangle mesh and a parameterized PSF emulating microscope optics. Gradients with respect to mesh and PSF parameters are computed via backpropagation, enabling their optimization to match real microscopy images.
3 Fluorescent 3D image rendering
Images can be represented as intensity maps $I$ from $[0,1]^{3}$ to $\mathbb{R}$ without loss of generality, as any non-cubic image can be linearly rescaled to fit within $[0,1]^{3}$. Following subsections introduce approximated models for both: imaging system and the geometry of biological samples, that are then integrated to construct our differentiable renderer.
Figure 3: Principle of 3D fluorescence microscopy rendering: A fluorophore distribution on the cell surface $\Lambda$ is convolved with the system’s PSF to produce a 3D fluorescence image.
@@ 3.1 Translation-invariant rendering model
We assume that the PSF $h$ of the fluorescence microscope is translation-invariant. Under this assumption, a smooth image $I:[0,1]^{3}\to\mathbb{R}$ is obtained by convolving a spatial fluorophore density $u_{\Lambda}:[0,1]^{3}\to\mathbb{R}$ with the point spread function (PSF) kernel $h:[0,1]^{3}\to\mathbb{R}$ :
$I(\textbf{x})=(u_{\Lambda}*h)(\textbf{x})=\int_{[0,1]^{3}}u_{\Lambda}(\textbf{p})h(\textbf{x}-\textbf{p})\mathrm{d}^{3}\textbf{p}.$ (1)
To circumvent the direct computation of these integrals, we perform the convolution in Fourier space by applying an element-wise multiplication of the Fourier transform (FT) $\hat{u}_{\Lambda}$ and $\hat{h}$, yielding: $\hat{I}=\hat{u}_{\Lambda}\cdot\hat{h}.$
With this image formation model, one can start from a first guess $\left(I^{0},h^{0}\right)$, and minimize the distance between the rendered image $I^{0}$ and a real microscopy image $I_{\text{GT}}$. Doing so allows to infer both the fluorophore density of the observed biological sample and the PSF of the imaging system: $\left(I^{*},h^{*}\right)$. However, to learn meaningful shape representations, both of these functions have to be parameterized using prior knowledge on the shape considered, its topology, as well as on the PSF.